### v2

# Sequence Reconstruction over the Deletion Channel

Abstract.
In this paper, we consider the Levenshtein’s sequence reconstruction problem in the case where the transmitted codeword is chosen from $\{0,1\}^{n}$ and the channel can delete up to $t$ symbols from the transmitted codeword. We determine the minimum number of channel outputs (assuming that they are distinct) required to reconstruct a list of size $\ell-1$ of candidate sequences, one of which corresponds to the original transmitted sequence. More specifically, we determine the maximum possible size of the intersection of $\ell\geqslant 3$ deletion balls of radius $t$ centered at $x_{1},x_{2},\dots,x_{\ell}$, where $x_{i}\in\{0,1\}^{n}$ for all $i\in\{1,2,\dots,\ell\}$ and $x_{i}\neq x_{j}$ for $i\neq j$, with $n\geqslant t+\ell-1$ and $t\geqslant 1$.
Institute for Systems Research and Department of ECE, University of Maryland, College Park, MD 20742, USA, fengxing@terpmail.umd.edu. Supported in part by NSF grant CCF 2330909.
1. Introduction
The study of sequence reconstruction was initiated by Levenshtein in [6, 5], where a sender transmits a codeword $x$ over multiple noisy channels. The receiver observes the outputs of these channels and attempts to uniquely reconstruct the transmitted codeword $x$. For the fixed codebook $\mathcal{C}$ and the channel the main task is to determine the minimum number of channel outputs required to guarantee unique reconstruction. The motivation for the sequence reconstruction problem originates from biology and chemistry, where traditional redundancy-based error correction methods are unsuitable. In recent years, the problem has regained attention due to its strong relevance to information retrieval in advanced storage technologies. In such systems, the stored data may consist of a single copy that is read multiple times or several redundant copies of the same information [4, 15]. This problem [4] is particularly significant in the context of DNA data storage [16, 2], where numerous noisy copies of DNA strands are available, and the objective is to accurately reconstruct the original information from these imperfect observations.
This sequence reconstruction problem has been extensively studied under various channels. Hirschberg and Regnier in [13] derived tight bounds on the number of string subsequences. In [6, 5], Levenshtein obtained the minimum number of channel outputs for the deletion and insertion channel required for unique reconstruction for the case where $\mathcal{C}$ consists of all binary vectors of length $n$. Gabrys and Yaakobi [3] later solved the sequence reconstruction problem over the $t$-deletion channel, where $\mathcal{C}$ consists of binary vectors such that $d_{L}(x,y)\geqslant 2$ for $x,y\in\mathcal{C}$, with $d_{L}(x,y)$ being the Levenshtein distance between $x$ and $y$. More recently, Pham, Goyal, and Kiah [11] obtained a complete asymptotic solution for this problem where $\mathcal{C}$ consists of binary vectors such that $d_{L}(x,y)\geqslant\ell$ for $x,y\in\mathcal{C}$.
The sequence reconstruction problem has also been studied under other channel models besides the deletion channel. Sala, Gabrys, Schoeny, and Dolecek [12] solved this problem for the insertion channel, assuming the codebook $\mathcal{C}$ consists of $q$-ary vectors satisfying $d_{L}(x,y)\geqslant\ell$ for all distinct $x,y\in\mathcal{C}$. Abu-Sini and Yaakobi [1] investigated the reconstruction problem for channels involving a single deletion combined with multiple substitutions, as well as channels involving a single insertion combined with a single substitution.
A variant of the sequence reconstruction problem allows the decoder to output a list of possible sequences instead of a unique reconstruction. Yaakobi and Bruck [14] studied this problem for channels introducing substitution errors. In particular, they investigated the maximum intersection of $m$ Hamming balls of radius $t$ centered at $x_{1},x_{2},\dots,x_{m}$, where $d_{H}(x_{i},x_{j})\geqslant d$ for $i\neq j$. Junnila, Laihonen, and Lehtilä [7, 8] analyzed the list size when the channel introduces substitution errors with $t=e+\ell$, where $e$ is the error-correcting capability of a binary code $\mathcal{C}$. More recently, they extended these results from the binary case to the $q$-ary case in [9].
In this paper we focus on the deletion channel. Formally, when a a codeword of length $n$ is sent through a $t$-deletion channel, a subsequence of length ntn-t is received. A $t$-deletion correcting code $\mathcal{C}$ is a subset of length-$n$ binary vectors such that for any vector $\in\mathcal{C}, xx$ can be uniquely identified from any length-(nt)(n-t) subsequence of $x$. More specifically, we study the minimum number of $t$-deletion channel outputs (assuming they are distinct) required to reconstruct a list of size $\ell-1$ of candidate sequences, one of which corresponds to the original transmitted sequence when $\mathcal{C}=\{0,1\}^{n}$. In other words, we determine the maximum possible size of the intersection of $\ell\geqslant 3$ deletion balls of radius $t$ centered at $x_{1},x_{2},\dots,x_{\ell}$, where $x_{i}\in\{0,1\}^{n}$ for all $i\in\{1,2,\dots,\ell\}$ and $x_{i}\neq x_{j}$ for $i\neq j$, with $n\geqslant t+\ell-1$ and $t\geqslant 1$.
2. Definitions and Preliminaries
Let $x$ be a binary sequence of length $n$ over $\mathbb{F}_{2}^{n}$. The deletion ball of radius $t$ centered at $x\in\mathbb{F}_{2}^{n}$ is define to be
$\mathcal{D}_{t}(x)=\{y\in\mathbb{F}_{2}^{n-t}|y\textrm{ is a subsequence of }x\}.$
For any two sequences $x_{1}$ and $x_{2}$, their Levenshtein distance is $t$ if $\mathcal{D}_{t}(x_{1})\cap\mathcal{D}_{t}(x_{2})\neq\emptyset$ and $\mathcal{D}_{t-1}(x_{1})\cap\mathcal{D}_{t-1}(x_{2})=\emptyset$
Let $a_{n}\in\mathbb{F}_{2}^{n}$ be an alternating sequences where its first bit is 1. For $0<t<n$, we denote the maximum size of a deletion ball of radius $t$, by $D(n,t)$,i.e
$D(n,t)=\max_{x\in\mathbb{F}_{2}^{n}}|\mathcal{D}_{t}(x)|.$
From [6] and [5], we know that
$D(n,t)=|\mathcal{D}_{t}(a_{n})|=\sum_{i=0}^{t}\binom{n-t}{i}$
and also
D(n,t)=D(n1,t)+D(n2,t1).D(n,t)=D(n-1,t)+D(n-2,t-1).
Note that $D(n,n)=1$ and $D(n,t)=0$ if $t<0$ or $n<t$.
Due to Levenshtein in [6] and [5], we have
$\max_{x_{1}\neq x_{2},x_{1},x_{2}\in\mathbb{F}_{2}^{n}}|\mathcal{D}_{t}(x_{1})\cap\mathcal{D}_{t}(x_{2})|=2D(n-2,t-1).$
In this paper we will study
$N(n,\ell,t):=\max_{x_{1}\neq x_{2}...\neq x_{\ell};x_{1},x_{2},...,x_{\ell}\in\mathbb{F}_{2}^{n}}|\mathcal{D}_{t}(x_{1})\cap\mathcal{D}_{t}(x_{2})...\cap\mathcal{D}_{t}(x_{\ell})|,$
where $\ell\geqslant 3$, $n\geqslant t+\ell-1$ and $t\geqslant 1$. Specifically, we establish the following theorem.
Theorem 1.
For $\ell\geqslant 3$, $n\geqslant t+\ell-1$ and $t\geqslant 1$, we have that
$N(n,\ell,t)=\sum_{i=1}^{\ell-2}D(n-2i,t-i)+2D(n-2(\ell-1),t-(\ell-1)).$
We will adopt the techniques and analysis from [3] and [11] to prove Theorem 1. In particular, in order to prove the upper bound in Theorem 1, we will use induction similar to that in [3] and [11]. Specifically, we will prove it by induction on $n,\ell,t$. The case where $\ell=3$, $t\geqslant 1$ and $n\geqslant t+\ell-1$ will serve as part of the base case. We could use the case $\ell=2$ as a base case but in the proof of the general case there are certain places where $\ell\geqslant 4$ is necessary and also we feel it is instructive to give the proof for $\ell=3$.
Before proceeding, we need to give some definitions and state some lemmas that will be used very often in our analysis.
Let $\chi\subset\mathbb{F}_{2}^{n}$ be a set and $v$ a sequence of length at most $n$. We denote by $\chi^{v}$ the set of all sequences in $\chi$ that start with the sequence $v$, that is,
$\chi^{v}=\{x\in\chi|v\textrm{ is a prefix of }x\}.$
For a sequence $v\in\mathbb{F}_{2}^{m}$ and a set $\chi\in\mathbb{F}_{2}^{n}$, the set $v\circ\chi$ is prepending the sequence $v$ before every sequence in $\chi$, that is
$v\circ\chi=\{(vx)|x\in\chi\}.$
The following two lemmas will be used very often in our analysis. Lemma 2 was derived in [3] and Lemma 3 was obtained in [10].
Lemma 2.
Let $n$, $m_{1}$, and $t$ be positive integers, and $x=x^{1}x^{2}\cdots x^{n}\in\mathbb{F}_{2}^{n}$, $x_{1}\in\mathbb{F}_{2}^{m_{1}}.$ Assume that $k$ is the smallest integer such that $x_{1}$ is a subsequence of $(x^{1},x^{2},...,x^{k})$. Then
$\mathcal{D}_{t}(x)^{x_{1}}=x_{1}\circ\mathcal{D}_{t^{*}}(x^{k+1},...,x^{n}),$
where $t^{*}=t-(k-m_{1}).$ In particular,
$|\mathcal{D}_{t}(x)^{x_{1}}|=|\mathcal{D}_{t^{*}}(x^{k+1},...,x^{n})|.$
Lemma 3.
Let $l<n$, $x\in\mathbb{F}_{2}^{n}$, $y\in\mathbb{F}_{2}^{n-l}$, where $y\in\mathcal{D}_{l}(x)$ and $l<t$. Then $\mathcal{D}_{t-l}(y)\subset\mathcal{D}_{t}(x)$.
For the ease of notation, let us define
$N_{\ell}(n,t):=\sum_{i=1}^{\ell-2}D(n-2i,t-i)+2D(n-2(\ell-1),t-(\ell-1)).$
3. The Intersection of Three Deletion Balls
In this section we will prove that
$N(n,3,t)=N_{3}(n,t).$
3.1. The Lower Bound
In this section we will show that 3D(n4,t2)+D(n3,t1)3D(n-4,t-2)+D(n-3,t-1) is a lower bound for $N(n,3,t)$ by showing that sequences $x_{1}=10a_{n-2}$, $x_{2}=01a_{n-2}$ and $x_{3}=0101a_{n-4}$ satisfy
$|\mathcal{D}_{t}(x_{1})\cap\mathcal{D}_{t}(x_{2})\cap\mathcal{D}_{t}(x_{3})|=3D(n-4,t-2)+D(n-3,t-1).$
We have the following Theorem.
Theorem 4.
For $t\geqslant 1$ and $n\geqslant t+2$,
$N(n,3,t)\geqslant 3D(n-4,t-2)+D(n-3,t-1).$
Proof.
Let $\chi=\mathcal{D}_{t}(x_{1})\cap\mathcal{D}_{t}(x_{2})\cap\mathcal{D}_{t}(x_{3})$. If $a=a^{1}a^{2}\cdots a^{j}$ is binary sequence of length $j$, then we denote $\bar{a}=\bar{a}^{1}\bar{a}^{2}\cdots\bar{a}^{j}$ as a sequence of length $j$ such that $\bar{a}^{i}=(1-a^{i})$ for $i\in\{1,2\cdots,j\}$. By Lemma 2 and Lemma 3, we have
$|\chi^{00}|$ $=|00\circ\mathcal{D}_{t-2}(a_{n-4})\cap\mathcal{D}_{t-2}(a_{n-4})\cap\mathcal{D}_{t-1}(1a_{n-4})|$
$=|\mathcal{D}_{t-2}(a_{n-4})|$
$=D(n-4,t-2).$
$|\chi^{01}|$ $=|01\circ\mathcal{D}_{t-1}(\bar{a}_{n-3})\cap\mathcal{D}_{t}(a_{n-2})\cap\mathcal{D}_{t}(01a_{n-4})|$
$=|\mathcal{D}_{t-1}(\bar{a}_{n-3})|$
$=D(n-3,t-1).$
$|\chi^{11}|$ $=|11\circ\mathcal{D}_{t-1}(\bar{a}_{n-3})\cap\mathcal{D}_{t-1}(\bar{a}_{n-3})\cap\mathcal{D}_{t-2}(a_{n-4})|$
$=|\mathcal{D}_{t-2}(a_{n-4})|$
$=D(n-4,t-2).$
$|\chi^{10}|$ $=|10\circ\mathcal{D}_{t}(a_{n-2})\cap\mathcal{D}_{t-2}(a_{n-4})\cap\mathcal{D}_{t-1}(1a_{n-4})|$
$=|\mathcal{D}_{t-2}(a_{n-4})|$
$=D(n-4,t-2).$
Since $\chi=\chi^{00}\cup\chi^{01}\cup\chi^{10}\cup\chi^{11}$, the proof is done. ∎
3.2. The Upper Bound
We will show that $N_{3}(n,t)$ is an upper bound for $N(n,3,t)$.
We will prove it by induction on $n$ and $t$. Let us first address the base case.
The base case is when $n=t+2$ and $t\geqslant 1$. since D(n2,t1)=2D(n-2,t-1)=2 and D(n4,t2)=1D(n-4,t-2)=1, $N_{3}(n,t)=4$. It is easy to see that after deleting $t$ symbols from each sequence of length $t+2$ there are only $2$ symbols left for each sequence. Since $N_{3}(n,t)=4$, $N(n,3,t)\leqslant N_{3}(n,t)$.
Now let $t=1$ and $n\geqslant t+2$. Then we have $N_{3}(n,t)=1$. By [5], we know that every two distinct binary sequences of length $n$ can have at most $2$ common supersequences of length $2$. Therefore, we have $N(n,3,1)\leqslant 1=N_{3}(n,1)$.
Now we will move onto the induction step. Assume that $N(n_{0},3,t_{0})\leqslant N_{3}(n_{0},t_{0})$ is true for all $n_{0}\geqslant t_{0}+2$ and $t_{0}\geqslant 1$ such that $n_{0}+t_{0}<n+t$. We will need a few lemmas to complete this step.
Lemma 5.
Assume that $t\geqslant 1$ and $n\geqslant t+2$. Let $x_{1}$$x_{2}$ and $x_{3}$ be three arbitrary sequences in $\mathbb{F}_{2}^{n}$ such that $x_{1}\neq x_{2}\neq x_{3}$ and $a=x_{1}^{1}=x_{2}^{1}=x_{3}^{1}$. Then
$|\mathcal{D}_{t}(x_{1})\cap\mathcal{D}_{t}(x_{2})\cap\mathcal{D}_{t}(x_{3})|\leqslant N_{3}(n,t).$
Proof.
Our proof follows the proof of Theorem 8 in [3]. Let $\chi=\mathcal{D}_{t}(x_{1})\cap\mathcal{D}_{t}(x_{2})\cap\mathcal{D}_{t}(x_{3}).$ We have
$|\chi^{a}|$ $=|a\circ\mathcal{D}_{t}(x_{1}^{2},...,x_{1}^{n})\cap\mathcal{D}_{t}(x_{2}^{2},...,x_{2}^{n})\cap\mathcal{D}_{t}(x_{3}^{2},...,x_{3}^{n}|$
$\leqslant\max_{x\neq y\neq z,x,y,z\in\mathbb{F}_{2}^{n-1}}|\mathcal{D}_{t}(x)\cap\mathcal{D}_{t}(y)\cap\mathcal{D}_{t}(z)|$
$\leqslant N_{3}(n-1,t).$
Suppose $x_{1}^{k}$ is the first occurrence of the symbol $\bar{a}$ in $x_{1}$ and the symbol $\bar{a}$ appears in $x_{1}$ not after it appears in $x_{2}$ and $x_{3}.$ If $x_{1}^{k}=x_{2}^{k}=x_{3}^{k}=\bar{a}$ we have
$|\chi^{\bar{a}}|$ $=|\bar{a}\circ\mathcal{D}_{t-(k-1)}(x_{1}^{k+1},...x_{1}^{n})\cap\mathcal{D}_{t-(k-1)}(x_{2}^{k+1},...,x_{2}^{n})\cap\mathcal{D}_{t-(k-1)}(x_{3}^{k+1},...,x_{3}^{n})|$
$\leqslant\max_{x\neq y\neq z,x,y,z\in\mathbb{F}_{2}^{n-k}}|\mathcal{D}_{t-(k-1)}(x)\cap\mathcal{D}_{t-(k-1)}(y)\cap\mathcal{D}_{t-(k-1)}(z)|$
$\leqslant N_{3}(n-k,t-k+1)$
$\leqslant N_{3}(n-2,t-1).$
If one of $x_{2}^{k}$ or $x_{3}^{k}$ is equal to $a$, say $x_{2}^{k}=a$ then we have
$|\chi^{\bar{a}}|$ $\leqslant|\bar{a}\circ\mathcal{D}_{t-k}(x_{2}^{k+2},...x_{2}^{n})|$
$\leqslant D(n-k-1,t-k)$
$\leqslant D(n-3,t-2)$
$\leqslant N_{3}(n-2,t-1).$
Since $|\chi^{a}|+|\chi^{\bar{a}}|\leqslant N_{3}(n-1,t)+N_{3}(n-2,t-1)=N_{3}(n,t)$, the proof is done. ∎
Due to Lemma  5, it is sufficient to consider any arbitrary sequences $x_{1}$, $x_{2}$ and $x_{3}$ in $\mathbb{F}_{2}^{n}$ such that $x_{1}^{1}=1$, $x_{2}^{1}=0$ and $x_{3}^{1}=0$ or $x_{1}^{1}=0$, $x_{2}^{1}=1$ and $x_{3}^{1}=1$. Due to the symmetry it suffices to deal with the case where $x_{1}^{1}=1$, $x_{2}^{1}=0$ and $x_{3}^{1}=0$. We need a few lemmas.
Lemma 6.
Assume that $t\geqslant 1$ and $n\geqslant t+2$. Let $x_{1}$$x_{2}$ and $x_{3}$ be three arbitrary sequences in $\mathbb{F}_{2}^{n}$ such that $x_{1}^{1}=1$, $x_{2}^{1}=0$ and $x_{3}^{1}=0$. If $x_{1}^{2}=x_{2}^{2}=x_{3}^{2}$, then
$|\mathcal{D}_{t}(x_{1})\cap\mathcal{D}_{t}(x_{2})\cap\mathcal{D}_{t}(x_{3})|\leqslant N_{3}(n,t).$
Proof.
Case 1: Assume that $x_{1}^{2}=x_{2}^{2}=x_{3}^{2}=0$.The we have $x_{1}=10x_{1}^{\prime}$, $x_{2}=00x_{2}^{\prime}$ and $x_{3}=00x_{3}^{\prime}$. Note that $x_{2}^{\prime}\neq x_{3}^{\prime}$. Suppose $x_{2}^{k}$ is the first occurrence of the symbol 1 in $x_{2}$ and the symbol 1 appears in $x_{2}$ not after it appears in $x_{3}$. If $x_{2}^{k}=x_{3}^{k}=1$ we have
$|\chi^{1}|$ $\leqslant|1\circ\mathcal{D}_{t-(k-1)}(x_{2}^{k+1},...,x_{2}^{n})\cap\mathcal{D}_{t-(k-1)}(x_{3}^{k+1},...,x_{3}^{n})|$
$\leqslant\max_{x\neq y,x,y\in\mathbb{F}_{2}^{n-k}}|\mathcal{D}_{t-(k-1)}(x)\cap\mathcal{D}_{t-(k-1)}(y)|$
$\leqslant 2D(n-2-k,t-k)$
$\leqslant 2D(n-5,t-3).$
If $x_{3}^{k}=0$ we have
$|\chi^{1}|$ $\leqslant|1\circ\mathcal{D}_{t-k}(x_{3}^{k+2},...,x_{3}^{n})|$
$\leqslant D(n-k-1,t-k)$
$\leqslant D(n-4,t-3)$
$\leqslant 2D(n-5,t-3).$
We also have that
$|\chi^{0}|$ $\leqslant|0\circ\mathcal{D}_{t-1}(x_{1}^{\prime})|$
$\leqslant D(n-2,t-1).$
Since D(n2,t1)=D(n3,t1)+D(n4,t2)D(n-2,t-1)=D(n-3,t-1)+D(n-4,t-2) and $D(n-5,t-3)\leqslant D(n-4,t-2)$, we have
$|\chi^{1}|+|\chi^{0}|\leqslant N_{3}(n,t).$
Case 2: Assume that $x_{1}^{2}=x_{2}^{2}=x_{3}^{2}=1$. Suppose $x_{1}^{k}$ is the first occurrence of the symbol 0 in $x_{1}$. We have that
$|\chi^{0}|$ $\leqslant|0\circ\mathcal{D}_{t-(k-1)}({x_{1}^{k+1},...,x_{1}^{n}})|$
$\leqslant D(n-k,t-k+1)$
$\leqslant D(n-3,t-2).$
We also have
$|\chi^{1}|$ $\leqslant|1\circ\mathcal{D}_{t-1}(\mathbf{x_{2}}^{\prime})\cap\mathcal{D}_{t-1}(\mathbf{x_{3}}^{\prime})|$
$\leqslant 2D(n-4,t-2).$
Since
$D(n-3,t-2)$ $=D(n-4,t-2)+D(n-5,t-3)$
$\leqslant D(n-3,t-1)+D(n-5,t-3),$
we have
$|\chi^{1}|+|\chi^{0}|\leqslant N_{3}(n,t).\qed$
Now we will show that it is sufficient to consider $x_{1}$$x_{2}$ and $x_{3}$ three arbitrary sequences in $\mathbb{F}_{2}^{n}$ such that $x_{1}^{1}=1$, $x_{2}^{1}=0$, $x_{3}^{1}=0$$x_{1}^{2}=0$, $x_{2}^{2}=1$ and $x_{3}^{2}=1$.
Lemma 7.
Assume $t\geqslant 1$ and $n\geqslant t+2$. Let $x_{1}$$x_{2}$ and $x_{3}$ be three arbitrary sequences in $\mathbb{F}_{2}^{n}$ such that $x_{1}^{1}=1$, $x_{2}^{1}=0$, $x_{3}^{1}=0$.
[…]
Then under (a),or (b), or (c)
$|\mathcal{D}_{t}(x_{1})\cap\mathcal{D}_{t}(x_{2})\cap\mathcal{D}_{t}(x_{3})|\leqslant N_{3}(n,t).$
Proof.
Let $\chi=\mathcal{D}_{t}(x_{1})\cap\mathcal{D}_{t}(x_{2})\cap\mathcal{D}_{t}(x_{3}).$
Under (a):
Let $x_{1}^{k}$ be the first occurrence of symbol 0 in $x_{1}$, we have
$|\chi^{0}|$ $\leqslant|0\circ\mathcal{D}_{t-(k-1)}(x_{1}^{k+1},...,x_{1}^{n})|$
$\leqslant D(n-k,t-k+1)$
$\leqslant D(n-3,t-2).$
Similarly, let $x_{2}^{k}$ be the first occurrence of symbol 1 in $x_{2}$, we have
$|\chi^{1}|$ $\leqslant|0\circ\mathcal{D}_{t-(k-1)}(x_{2}^{k+1},...,x_{2}^{n})|$
$\leqslant D(n-k,t-k+1)$
$\leqslant D(n-3,t-2).$
Therefore, we have
$|\chi^{1}|+|\chi^{0}|\leqslant N_{3}(n,t).$
Under (b):
Let $x_{1}^{k}$ be the first occurrence of symbol 0 in $x_{1}$ we have
$|\chi^{0}|$ $\leqslant|0\circ\mathcal{D}_{t-(k-1)}(x_{1}^{k+1},...,x_{1}^{n})|$
$\leqslant D(n-k,t-k+1)$
$\leqslant D(n-3,t-2).$
Similarly, let $x_{3}^{k}$ be the first occurrence of symbol 1 in $x_{3}$, we have
$|\chi^{1}|$ $\leqslant|0\circ\mathcal{D}_{t-(k-1)}(x_{3}^{k+1},...,x_{3}^{n})|$
$\leqslant D(n-k,t-k+1)$
$\leqslant D(n-3,t-2).$
Therefore, we have
$|\chi^{1}|+|\chi^{0}|\leqslant N_{3}(n,t).$
Under(c):
If $x_{3}^{k}$ be the first occurrence of symbol 1 in $x_{3}$, we have
$|\chi^{1}|$ $\leqslant|0\circ\mathcal{D}_{t-(k-1)}(x_{3}^{k+1},...,x_{3}^{n})|$
$\leqslant D(n-k,t-k+1)$
$\leqslant D(n-3,t-2).$
$|\chi^{0}|$ $\leqslant|0\circ\mathcal{D}_{(t-1)}(x_{1}^{3},...,x_{1}^{n})|$
$\leqslant D(n-2,t-1).$
Therefore, we have
$|\chi^{1}|+|\chi^{0}|\leqslant N_{3}(n,t).\qed$
Lemma 8.
Assume that $t\geqslant 1$ and $n\geqslant t+2$. Let $x_{1}$$x_{2}$ and $x_{3}$ be three arbitrary sequences in $\mathbb{F}_{2}^{n}$ such that $x_{1}^{1}=1$, $x_{2}^{1}=0$ and $x_{3}^{1}=0$ and $x_{1}^{2}=0$, $x_{2}^{2}=1$ and $x_{3}^{1}=1$ Then we have
$|\mathcal{D}_{t}(x_{1})\cap\mathcal{D}_{t}(x_{2})\cap\mathcal{D}_{t}(x_{3})|\leqslant N_{3}(n,t).$
Proof.
Let $\chi=\mathcal{D}_{t}(x_{1})\cap\mathcal{D}_{t}(x_{2})\cap\mathcal{D}_{t}(x_{3}).$ We have
$|\chi^{1}|$ $\leqslant|1\circ\mathcal{D}_{t-1}(x_{2}^{3},...,x_{2}^{n})\cap\mathcal{D}_{t-1}(x_{3}^{3},...,x_{3}^{n})|$
$\leqslant\max_{x\neq y,x,y\in\mathbb{F}_{2}^{n-2}}|\mathcal{D}_{t-1}(x)\cap\mathcal{D}_{t-1}(y)|$
$\leqslant 2D(n-4,t-2).$
Also
$|\chi^{0}|$ $\leqslant|0\circ\mathcal{D}_{(t-1)}(x_{1}^{3},...,x_{1}^{n})|$
$\leqslant D(n-2,t-1).$
Since $2D(n-4,t-2)+D(n-2,t-1)=N_{3}(n,t)$, the proof is done. ∎
By Lemmas 5, 6, 7, and 8, the induction step is finished and we have the following theorem.
Theorem 9.
For $t\geqslant 1$ and $n\geqslant t+2$, we have
$N(n,3,t)\leqslant N_{3}(n,t)$
4. The Intersection of $\ell\geqslant 4$ Deletion Balls
In this section we will prove that
$N(n,\ell,t)=N_{\ell}(n,t):=\sum_{i=1}^{\ell-2}D(n-2i,t-i)+2D(n-2(\ell-1),t-(\ell-1)).$
4.1. The Lower Bound
We will show that $N_{\ell}(n,t)$ is a lower bound for $N(n,\ell,t)$ by showing that $x_{1}=10a_{n-2}$, $x_{2}=01a_{n-2}$, $x_{j}=\underbrace{0101\cdots 01}_{2(j-1)}a_{n-2(j-1)}$ for $j\in\{2,\cdots,\ell\}$ satisfy
$|\cap_{i=1}^{\ell}\mathcal{D}_{t}(x_{i})|=N_{\ell}(n,t).$
Theorem 10.
For $\ell\geqslant 4$, $t\geqslant 1$, and $n\geqslant t+\ell-1$,
$N(n,\ell,t)\geqslant\sum_{i=1}^{\ell-2}D(n-2i,t-i)+2D(n-2(\ell-1),t-(\ell-1))$
Proof.
Let $x_{1}=10a_{n-2}$, $x_{2}=01a_{n-2}$, $x_{j}=\underbrace{0101\cdots 01}_{2(j-1)}a_{n-2(j-1)}$ for $j\in\{2,\cdots,\ell\}$ and let $\chi=\cap_{i=1}^{\ell}\mathcal{D}_{t}(x_{i})$. Observe that
$(0\circ\mathcal{D}_{(t-1)}(x_{1}^{3},...,x_{1}^{n}))\cap(0\circ\mathcal{D}_{t}(x_{2}^{2},...,x_{2}^{n}))\cap\cdots\cap(0\circ\mathcal{D}_{t}(x_{\ell}^{2},...,x_{\ell}^{n}))=0\circ\mathcal{D}_{(t-1)}(x_{1}^{3},...,x_{1}^{n})$
and
$(1\circ\mathcal{D}_{t}(x_{1}^{2},...,x_{1}^{n}))\cap(1\circ\mathcal{D}_{t-1}(x_{2}^{3},...,x_{2}^{n}))\cap\cdots\cap(1\circ\mathcal{D}_{t-1}(x_{\ell}^{3},...,x_{\ell}^{n}))$
$=(1\circ\mathcal{D}_{t-1}(x_{2}^{3},...,x_{2}^{n}))\cap(1\circ\mathcal{D}_{t-1}(x_{3}^{3},...,x_{3}^{n}))\cdots\cap(1\circ\mathcal{D}_{t-1}(x_{\ell}^{3},...,x_{\ell}^{n}))$
By Lemma 1 and Lemma 2, we have
$|\chi^{0}|$ $=|(0\circ\mathcal{D}_{(t-1)}(x_{1}^{3},...,x_{1}^{n}))\cap(0\circ\mathcal{D}_{t}(x_{2}^{2},...,x_{2}^{n}))\cap\cdots\cap(0\circ\mathcal{D}_{t}(x_{\ell}^{2},...,x_{\ell}^{n}))|$
$=|(0\circ\mathcal{D}_{t-1}(x_{1}^{3},...,x_{1}^{n}))|$
$=D(n-2,t-1)$
$|\chi^{1}|$ $=|(1\circ\mathcal{D}_{t}(x_{1}^{2},...,x_{1}^{n}))\cap(1\circ\mathcal{D}_{t-1}(x_{2}^{3},...,x_{2}^{n}))\cap(1\circ\mathcal{D}_{t-1}(x_{3}^{3},...,x_{3}^{n}))\cap\cdots\cap(1\circ\mathcal{D}_{t-1}(x_{\ell}^{3},...,x_{\ell}^{n}))|$
$=|(1\circ\mathcal{D}_{t-1}(x_{2}^{3},...,x_{2}^{n}))\cap(1\circ\mathcal{D}_{t-1}(x_{3}^{3},...,x_{3}^{n}))\cdots\cap(1\circ\mathcal{D}_{t-1}(x_{\ell}^{3},...,x_{\ell}^{n}))|$
$=N_{\ell-1}(n-2,t-1)$
Therefore,
$|\chi|$ $=|\chi^{0}|+|\chi^{1}|$
$=D(n-2,t-1)+N_{\ell-1}(n-2,t-1).\qed$
Now we have the following recursive relations.
$N_{2}(n,t)=2D(n-2,t-1)$
$N_{3}(n,t)=D(n-2,t-1)+2D(n-4,t-2)$
$N_{4}(n,t)=D(n-2,t-1)+N_{3}(n-2,t-1)$
$N_{\ell}(n,t)=D(n-2,t-1)+N_{\ell-1}(n-2,t-1).$
Therefore,
$N_{\ell}(n,t)=\sum_{i=1}^{\ell-2}D(n-2i,t-i)+2D(n-2(\ell-1),t-(\ell-1)).$
4.2. The Upper Bound
We will show that $N_{\ell}(n,t)$ is an upper bound for $N(n,\ell,t)$ by induction on $\ell$, $n$, $t$. Let us first address the base case.
The base case is: $n=t+\ell-1$, $t\geqslant 1$, and $\ell\geqslant 3$ or
$t=1$, $\ell\geqslant 3$, and $n\geqslant t+\ell-1$ or
$\ell=3$ , $t\geqslant 1$, and $n\geqslant t+\ell-1$.
$n=t+\ell-1$, $t\geqslant 1$, and $\ell\geqslant 3$, it is easy to calculate that $N_{\ell}(t+\ell-1,t)=2^{\ell-1}$. After deleting $t$ symbols from each sequence of length $t+\ell-1$ there are only $\ell-1$ symbols left for each sequence. Since $N_{\ell}(n,t)=2^{\ell-1}$, $N(n,\ell,t)\leqslant N_{\ell}(n,t)$.
For $t=1$, $\ell\geqslant 3$, and $n\geqslant t+\ell-1$, we have $N_{\ell}(n,1)=1$. Since $N(n,3,1)=1$, we have $N(n,\ell,1)\leqslant 1=N_{\ell}(n,1)$.
For $\ell=3$ , $t\geqslant 1$, and $n\geqslant t+\ell-1$, in the previous section it has been proven.
Now we will move onto the induction step. Assume that $N(n_{0},\ell_{0},t_{0})\leqslant N_{\ell_{0}}(n_{0},t_{0})$ is true for all $n_{0}\geqslant t_{0}+2$, $\ell_{0}\geqslant 3$, and $t_{0}\geqslant 1$ such that $n_{0}+t_{0}+\ell_{0}<n+t+\ell$. We will need a few lemmas to complete this step.
Let us first show the following two lemmas which will be used in our analysis several times.
Lemma 11.
For any integers $x,y$ and any $m\geqslant 2$,
$D(x-1,y-1)\ \leqslant\ N_{m}(x,y).$
Proof.
We argue by induction on $m\geqslant 2$.
For the base case let $m=2$. Here $N_{2}(x,y)=2D(x-2,y-1)$.
We have
D(x1,y1)=D(x2,y1)+D(x3,y2).D(x-1,y-1)=D(x-2,y-1)+D(x-3,y-2).
Moreover,
$D(x-3,y-2)\ \leqslant\ D(x-2,y-1),$
Thus
[…]
For the induction step assume the claim holds for some $m-1\geqslant 2$, i.e.,
$D(u-1,v-1)\ \leqslant\ N_{m-1}(u,v)\qquad\text{for all }u,v\in\mathbb{Z}.$
We show it for $m$. Using the same recursion,
D(x1,y1)=D(x2,y1)+D(x3,y2).D(x-1,y-1)=D(x-2,y-1)+D(x-3,y-2).
On the other hand, by the definition of $N_{m}$,
$N_{m}(x,y)=D(x-2,y-1)+N_{m-1}(x-2,y-1).$
Applying the induction hypothesis gives
$D(x-3,y-2)\ \leqslant\ N_{m-1}(x-2,y-1).$
Therefore
$D(x-1,y-1)\ \leqslant\ D(x-2,y-1)+N_{m-1}(x-2,y-1)\ =\ N_{m}(x,y),$
completing the induction. ∎
Lemma 12.
For any $m$ and any $L\geqslant m$,
$N_{m}(u-1,v-1)\ \leqslant\ N_{L}(u,v).$
Proof.
Let $L=m$. Since $D(n,t)\geqslant D(n-1,t-1)$, we have
$N_{m}(u-1,v-1)\leqslant N_{L}(u,v).$
Now assume that $L>m$. We will bound $N_{m}(u-1,v-1)$ from above termwise.
For $1\leqslant i\leqslant m-2$, we have:
$D(u-1-2i,\,v-1-i)\ \leqslant\ D(u-2i,\,v-i).$
Since $L>m$, the indices $i=1,\dots,m-2$ all lie in the index range $1\leqslant i\leqslant L-2$ of $N_{L}(u,v)$. Thus each of these terms is dominated by a corresponding term of $N_{L}(u,v)$.
Let a=u12(m1)a=u-1-2(m-1) and b=v1(m1)b=v-1-(m-1). Then
$2D(a,b)=2D\bigl(u-1-2(m-1),\,v-1-(m-1)\bigr).$
If $L>m$, then $m-1\leqslant L-2$. We have
$2D(a,b)\leqslant D\bigl(u-2(m-1)+1,\,v-(m-1)\bigr)\leqslant D\bigl(u-2(m-1),\,v-(m-1)\bigr).$
Therefore, we have
$N_{m}(u-1,v-1)\ \leqslant\ N_{L}(u,v).\qed$
Now let us assume that $n\geqslant t+\ell-1$, $t\geqslant 1$ and $\ell\geqslant 4$.
Lemma 13.
Suppose that among the vectors $x_{1},x_{2},\dots,x_{\ell}$, $\ell_{1}$ of them begin with $10$ and $\ell_{2}$ of them begin with $01$, where $\ell_{1}\geqslant 2$ and $\ell_{2}\geqslant 2$. Then
$|\mathcal{D}_{t}(\mathbf{x}_{1})\cap\mathcal{D}_{t}(\mathbf{x}_{2})\cap...\cap\mathcal{D}_{t}(\mathbf{x}_{\ell})|\leqslant N_{\ell}(n,t).$
Proof.
Let $\chi=\cap_{i=1}^{\ell}\mathcal{D}_{t}(x_{i})$ and then $\chi^{0}=0\circ\cap_{i=1}^{\ell_{1}}\mathcal{D}_{t-1}(x_{i}^{3}x_{i}^{4}\cdots x_{i}^{n})$.
$|\chi^{0}|\leqslant\max_{y_{i}\in\{0,1\}^{n-2}}\cap_{i=1}^{\ell_{1}}\mathcal{D}_{t-1}(y_{i})\leqslant N_{\ell_{1}}(n-2,t-1).$
Similarly, we have
$|\chi^{1}|\leqslant\max_{y_{i}\in\{0,1\}^{n-2}}\cap_{i=1}^{\ell_{2}}\mathcal{D}_{t-1}(y_{i})\leqslant N_{\ell_{2}}(n-2,t-1).$
Let $\ell,\ell_{1},\ell_{2}\geqslant 2$ with $\ell_{1}+\ell_{2}=\ell$ and $\ell\geqslant 4$. We will show that
$N_{\ell}(n,t)\geqslant N_{\ell_{1}}(n-2,t-1)+N_{\ell_{2}}(n-2,t-1).$
We first note the identity
D(n,t)=D(n1,t)+D(n2,t1).D(n,t)=D(n-1,t)+D(n-2,t-1). (1)
Since
$D(n-1,t)-D(n-2,t-1)=\binom{n-t-1}{t}\geqslant 0,$
we obtain from (1) that
$D(n,t)\geqslant 2D(n-2,t-1).$ (2)
Applying (2) to each term in the definition of $N_{\ell}(n,t)$ gives
$N_{\ell}(n,t)$ $=\sum_{i=1}^{\ell-2}D(n-2i,t-i)\;+\;2D(n-2(\ell-1),t-(\ell-1))$ (3)
$\geqslant 2\sum_{i=1}^{\ell-2}D(n-2-2i,t-1-i)\;+\;2D(n-2-2(\ell-1),t-\ell).$ (4)
Now we expand $N_{\ell_{1}}(n-2,t-1)$ and $N_{\ell_{2}}(n-2,t-1)$:
$N_{\ell_{1}}(n-2,t-1)$ $=\sum_{i=1}^{\ell_{1}-2}D(n-2-2i,t-1-i)\;+\;2D(n-2-2(\ell_{1}-1),t-\ell_{1}),$
$N_{\ell_{2}}(n-2,t-1)$ $=\sum_{j=1}^{\ell_{2}-2}D(n-2-2j,t-1-j)\;+\;2D(n-2-2(\ell_{2}-1),t-\ell_{2}).$
Summing, we obtain
$N_{\ell_{1}}(n-2,t-1)+N_{\ell_{2}}(n-2,t-1)$ (5)
$=\sum_{i=1}^{\ell_{1}-2}D(n-2-2i,t-1-i)+\sum_{j=1}^{\ell_{2}-2}D(n-2-2j,t-1-j)$ (6)
$+2D(n-2-2(\ell_{1}-1),t-\ell_{1})+2D(n-2-2(\ell_{2}-1),t-\ell_{2}).$ (7)
Also we have
$D(n-2(\ell_{1}-1),\,t-(\ell_{1}-1))\geqslant 2D(n-2-2(\ell_{1}-1),\,t-\ell_{1}),$
and
$D(n-2(\ell_{2}-1),\,t-(\ell_{2}-1))\geqslant 2D(n-2-2(\ell_{2}-1),\,t-\ell_{2}).$
Since $\ell_{1}-1,\ell_{2}-1\leqslant\ell-2$ when $\ell_{1},\ell_{2}\geqslant 2$, every term in (7) appears among the summands in (4) with coefficient at most $2$. The additional term $2D(n-2-2(\ell-1),t-\ell)\geqslant 0$ in (4) only increases the right-hand side of that inequality.
Therefore, comparing (4) and (7) yields
$N_{\ell}(n,t)\geqslant N_{\ell_{1}}(n-2,t-1)+N_{\ell_{2}}(n-2,t-1),$
which completes the proof. ∎
Lemma 14.
Let $x_{1}$$x_{2}$$\cdots$ and $x_{\ell}$ be $\ell$ arbitrary sequences in $\mathbb{F}_{2}^{n}$ such that $x_{i}\neq x_{j}$ for $i\neq j$ and $a=x_{1}^{1}=x_{2}^{1}=\cdots=x_{\ell}^{1}$. Then
$|\cap_{i}^{\ell}\mathcal{D}_{t}(x_{i})|\leqslant N_{\ell}(n,t).$
Proof.
Our proof follows the proof of Theorem 8 in [3]. Let $\chi=\mathcal{D}_{t}(x_{1})\cap\mathcal{D}_{t}(x_{2})\cap\cdots\cap\mathcal{D}_{t}(x_{\ell})$. We have
$|\chi^{a}|$ $=|a\circ\mathcal{D}_{t}(x_{1}^{2},...,x_{1}^{n})\cap\mathcal{D}_{t}(x_{2}^{2},...,x_{2}^{n})\cap\cdots\cap\mathcal{D}_{t}(x_{\ell}^{2},...,x_{\ell}^{n})|$
$\leqslant\max_{x_{1}\neq x_{2}\neq\cdots x_{\ell};x_{1},x_{2},\cdots,x_{\ell}\in\mathbb{F}_{2}^{n-1}}|\mathcal{D}_{t}(x_{1})\cap\mathcal{D}_{t}(x_{2})\cap\cdots\cap\mathcal{D}_{t}(x_{\ell})|$
$\leqslant N_{\ell}(n-1,t)$
Suppose $x_{1}^{k}$ is the first occurrence of the symbol $\bar{a}$ in $x_{1}$ and the symbol $\bar{a}$ appears in $x_{1}$ not after it appears in $x_{2}$, $x_{3}$, $\cdots$ and $x_{\ell}$.
If $x_{1}^{k}=x_{2}^{k}=\cdots=x_{\ell}^{k}=\bar{a}$ we have
$|\chi^{\bar{a}}|$ $=|\bar{a}\circ\mathcal{D}_{t-(k-1)}(x_{1}^{k+1},...x_{1}^{n})\cap\mathcal{D}_{t-(k-1)}(x_{2}^{k+1},...,x_{2}^{n})\cap\cdots\cap\mathcal{D}_{t-(k-1)}(x_{\ell}^{k+1},...,x_{\ell}^{n})|$
$\leqslant\max_{\mathbf{x_{1}}\neq\mathbf{x_{2}}\cdots\neq\mathbf{x_{\ell}};x_{1},x_{2},\cdots,x_{\ell}\in\mathbb{F}_{2}^{n-k}}|\mathcal{D}_{t-(k-1)}(x_{1})\cap\mathcal{D}_{t-(k-1)}(x_{2})\cap\cdots\cap\mathcal{D}_{t-(k-1)}(x_{\ell})|$
$\leqslant N_{\ell}(n-k,t-k+1)$
$\leqslant N_{\ell}(n-2,t-1)$
If one of $x_{2}^{k}$, $x_{3}^{k}$ $\cdots$ or $x_{\ell}^{k}$ is equal to $a$, say $x_{2}^{k}=a$ then we have
$|\chi^{\bar{a}}|$ $\leqslant|\bar{a}\circ\mathcal{D}_{t-k}(x_{2}^{k+2},...x_{2}^{n})|$ (8)
$\leqslant D(n-k-1,t-k)$ (9)
$\leqslant D(n-3,t-2)$ (10)
$\leqslant N_{\ell}(n-2,t-1)$ (11)
Let us show (11).
By using (1), we prove by induction that for any integer $k\geqslant 1$,
$D(n-3,t-2)=\sum_{i=1}^{k}D(n-2-2i,\ t-1-i)+D(n-(2k+3),\ t-(k+2)).$ (12)
Let $k=\ell-2$. Then (12) becomes
$D(n-3,t-2)=\sum_{i=1}^{\ell-2}D(n-2-2i,\ t-1-i)+D\bigl(n-(2\ell-1),\ t-\ell).$ (13)
Note that from
D(n,t)=D(n1,t)+D(n2,t1),D(n,t)=D(n-1,t)+D(n-2,t-1), (14)
and
$D(n-1,t)-D(n-2,t-1)=\binom{n-t-1}{t}\geqslant 0,$
we have $D(n,t)\leqslant 2D(n-1,t)$. Thus
$D(n-(2\ell-1),\,t-\ell)\leqslant 2\,D(n-2\ell,\,t-\ell)=2\,D(n-2-2(\ell-1),\,t-\ell).$
Substituting this estimate into (13) yields
$D(n-3,t-2)\leqslant\sum_{i=1}^{\ell-2}D(n-2-2i,\ t-1-i)+2\,D(n-2-2(\ell-1),\,t-\ell)=N_{\ell}(n-2,t-1).$
Since $|\chi^{a}|+|\chi^{\bar{a}}|\leqslant N_{\ell}(n-1,t)+N_{\ell}(n-2,t-1)=N_{\ell}(n,t)$, the proof is completed. ∎
Lemma 15.
Suppose that among the vectors $x_{1},x_{2},\dots,x_{\ell}$, $\ell_{1}\geqslant 1$ of them begin with $11$ and $\ell_{3}\geqslant 1$ of them begin with $00$. Then
$|\mathcal{D}_{t}(x_{1})\cap\mathcal{D}_{t}(x_{2})\cap...\cap\mathcal{D}_{t}(x_{\ell})|\leqslant N_{\ell}(n,t).$
Proof.
Assume that the first two coordinates of $x_{1}$, $x_{2}$$\cdots,$ $x_{\ell_{1}}$ are $11$ and the first two coordinates of $x_{\ell_{1}+1}$, $x_{\ell_{1}+2}$, $\cdots$, $x_{\ell_{1}+\ell_{3}}$ are $00$. Let $\chi=\cap_{i=1}^{\ell}\mathcal{D}_{t}(x_{i})$
Case 1: $\ell_{1}\geqslant 2$ and $\ell_{3}\geqslant 2$.
Then we have $\chi^{0}\subset 0\circ\max_{y_{i}:y_{i}\in\{0,1\}^{n-3}}\cap_{i=1}^{\ell_{1}}\mathcal{D}_{t-2}(y_{i})$.
$|\chi^{0}|\leqslant\max_{y_{i}\in\{0,1\}^{n-3},y_{i}\neq y_{j}}\cap_{i=1}^{\ell_{1}}\mathcal{D}_{t-2}(y_{i})\leqslant N_{\ell_{1}}(n-3,t-2).$
Similarly, we have $\chi^{1}\subset 1\circ\max_{y_{i}:y_{i}\in\{0,1\}^{n-3}}\cap_{i=1}^{\ell_{3}}\mathcal{D}_{t-2}(y_{i})$.
$|\chi^{0}|\leqslant\max_{y_{i}\in\{0,1\}^{n-3},y_{i}\neq y_{j}}\cap_{i=1}^{\ell_{3}}\mathcal{D}_{t-2}(y_{i})\leqslant N_{\ell_{3}}(n-3,t-2).$
By the proof of Lemma  13, we have
$N_{\ell_{1}}(n-2,t-1)+N_{\ell_{3}}(n-2,t-1)\leqslant\ N_{\ell}(n,t).$
Therefore,
$N_{\ell_{1}}(n-3,t-2)+N_{\ell_{3}}(n-3,t-2)\ \leqslant\ N_{\ell_{1}}(n-2,t-1)+N_{\ell_{3}}(n-2,t-1)\ \leqslant\ N_{\ell}(n,t).$
Case 2: $\ell_{1}=1$ and $\ell_{3}\geqslant 2$. $\chi^{0}\subset 0\circ\max_{y:y\in\{0,1\}^{n-3}}\mathcal{D}_{t-2}(y)$.
$|\chi^{0}|\leqslant\max_{y_{i}\in\{0,1\}^{n-3}}\mathcal{D}_{t-2}(y_{i})\leqslant D(n-3,t-2).$
Similarly, we have $\chi^{1}\subset 1\circ\max_{y_{i}:y_{i}\in\{0,1\}^{n-3}}\cap_{i=1}^{\ell_{3}}\mathcal{D}_{t-2}(y_{i})$.
$|\chi^{0}|\leqslant\max_{y_{i}\in\{0,1\}^{n-3},y_{i}\neq y_{j}}\cap_{i=1}^{\ell_{3}}\mathcal{D}_{t-2}(y_{i})\leqslant N_{\ell_{3}}(n-3,t-2).$
We claim that $D(n-3,t-2)+N_{\ell_{3}}(n-3,t-2)\leqslant N_{\ell}(n,t)$. Let us prove this claim.
We have
$N_{\ell}(n,t)=D(n-2,t-1)\;+\;N_{\ell-1}(n-2,t-1).$
and
$D(n-3,t-2)\ \leqslant\ D(n-2,t-1).$
Apply Lemma  12 with $m=\ell_{3}$, $L=\ell-1$, and (u,v)=(n2,t1)(u,v)=(n-2,t-1) to get
$N_{\ell_{3}}(n-3,t-2)\ \leqslant\ N_{\ell-1}(n-2,t-1),$
since $\ell_{3}\leqslant\ell-1$ by hypothesis.
Therefore,
$D(n-3,t-2)+N_{\ell_{3}}(n-3,t-2)\ \leqslant\ D(n-2,t-1)+N_{\ell-1}(n-2,t-1)\ =\ N_{\ell}(n,t).$
Case 3: $\ell_{1}=1$ and $\ell_{3}=1$. We have $\chi^{0}\subset 0\circ\max_{y:y\in\{0,1\}^{n-3}}\mathcal{D}_{t-2}(y)$.
$|\chi^{0}|\leqslant\max_{y_{i}\in\{0,1\}^{n-3}}\mathcal{D}_{t-2}(y_{i})\leqslant D(n-3,t-2),$
and $\chi^{1}\subset 1\circ\max_{y:y\in\{0,1\}^{n-3}}\mathcal{D}_{t-2}(y)$.
$|\chi^{0}|\leqslant\max_{y_{i}\in\{0,1\}^{n-3}}\mathcal{D}_{t-2}(y_{i})\leqslant D(n-3,t-2).$
By Lemma  11, we have
$2D(n-3,t-2)\leqslant 2N_{2}(n-2,t-1)\leqslant N_{\ell}(n,t),$
since $\ell\geqslant 4$. ∎
Now we can assume none of the $\ell$ vectors start with $00$.
Lemma 16.
Suppose that among the vectors $x_{1},x_{2},\dots,x_{\ell}$, $\ell_{1}\geqslant 1$ of them begin with $11$, $\ell_{2}\geqslant 1$ of them begin with $01$, and $\ell_{4}$ of them begin with $10$. Then
$|\mathcal{D}_{t}(x_{1})\cap\mathcal{D}_{t}(x_{2})\cap...\cap\mathcal{D}_{t}(x_{\ell})|\leqslant N_{\ell}(n,t).$
Proof.
Let $\chi=\cap_{i=1}^{\ell}\mathcal{D}_{t}(x_{i})$
Case 1: $\ell_{1}\geqslant 1$ and $\ell_{2}\geqslant 2$.
Assume that the first two coordinates of $x_{1}$$x_{2}$$\cdots,$ $x_{\ell_{1}}$ are $11$ and the first two coordinates of $x_{\ell_{1}+1}$$x_{\ell_{1}+2}$$\cdots,$ $x_{\ell_{1}+\ell_{2}}$ are $01$. Then we have $\chi^{0}\subset 0\circ\max_{y:y\in\{0,1\}^{n-3}}\mathcal{D}_{t-2}(y)$.
$|\chi^{0}|\leqslant\max_{y_{i}\in\{0,1\}^{n-3}}\mathcal{D}_{t-2}(y_{i})\leqslant D(n-3,t-2).$
Similarly, $\chi^{1}\subset 1\circ\max_{y_{1},y_{2}\in\{0,1\}^{n-2}}\mathcal{D}_{t-1}(y_{1})\cap\mathcal{D}_{t-1}(y_{2})$.
$|\chi^{1}|\leqslant\max_{y_{1},y_{2}\in\{0,1\}^{n-2}}\mathcal{D}_{t-1}(y_{1})\cap\mathcal{D}_{t-1}(y_{2})\leqslant 2D(n-4,t-2).$
We have
$N_{\ell}(n,t)$ $=D(n-2,t-1)+N_{\ell-1}(n-2,t-1).$ (15)
Note that we have
D(n2,t1) 2D(n4,t2).D(n-2,t-1)\ $\geqslant\$2\,D(n-4,t-2).
By Lemma  11:
$N_{\ell-1}(n-2,t-1)\ \geqslant\ D\bigl((n-2)-1,(t-1)-1\bigr)\ =\ D(n-3,t-2).$
Combining these two bounds in (15) yields
$N_{\ell}(n,t)\ \geqslant\ 2\,D(n-4,t-2)\;+\;D(n-3,t-2),$
as claimed.
Case 2: $\ell_{1}\geqslant 2$ and $\ell_{2}=1$.
We have $\chi^{1}\subset 1\circ\max_{y\in\{0,1\}^{n-2}}\mathcal{D}_{t-1}(y)$.
$|\chi^{1}|\leqslant\max_{y\in\{0,1\}^{n-2}}\mathcal{D}_{t-1}(y)\leqslant D(n-2,t-1).$
$\chi^{0}\subset 0\circ\max_{y_{i}:y_{i}\in\{0,1\}^{n-3}}\cap_{i=1}^{\ell_{1}}\mathcal{D}_{t-2}(y_{i})$.
$|\chi^{0}|\leqslant\max_{y_{i}\in\{0,1\}^{n-3},y_{i}\neq y_{j}}\cap_{i=1}^{\ell_{1}}\mathcal{D}_{t-2}(y_{i})\leqslant N_{\ell_{1}}(n-3,t-2).$
By lemma  11, we have
$D(n-2,t-1)\leqslant N_{\ell}(n-1,t).$
Additionally we have $N_{\ell_{1}}(n-3,t-2)\leqslant N_{\ell}(n-2,t-1)$ by Lemma  12.
Finally, $N_{\ell}(n-2,t-1)+N_{\ell}(n-1,t)=N_{\ell}(n,t)$
Case 3: $\ell_{1}=1$, $\ell_{2}=1$ and $\ell_{4}=\ell-2$.
We have $\chi^{1}\subset 1\circ\max_{y\in\{0,1\}^{n-2}}\mathcal{D}_{t-1}(y)$.
$|\chi^{1}|\leqslant\max_{y\in\{0,1\}^{n-2}}\mathcal{D}_{t-1}(y)\leqslant D(n-2,t-1).$
Also we have since $k\geqslant 3$
$\chi^{0}$ $\subset 0\circ\max_{x_{i}\in\{0,1\}^{n-2},y\in\{0,1\}^{n-k}}\cap_{i=1}^{\ell_{4}}\mathcal{D}_{t-1}(x_{i})\cap\mathcal{D}_{t-k+1}(y)$
$\subset 0\circ\max_{y\in\{0,1\}^{n-3}}\mathcal{D}_{t-2}(y).$
we have $|\chi^{0}|\leqslant D(n-3,t-2)\leqslant N_{\ell}(n-2,t-1).$ By Lemma  11, we have $D(n-2,t-1)\leqslant N_{\ell}(n-1,t)$. ∎
Suppose that among the vectors $x_{1},x_{2},\dots,x_{\ell}$, one begins with $10$ and another begins with $11$. In this case, by symmetry, the situation is equivalent to that considered in the preceding lemma.
Now the induction step is finished. we have the following theorem.
Theorem 17.
For $t\geqslant 1$, $\ell\geqslant 3$, and $n\geqslant t+\ell-1$, we have
$N(n,\ell,t)\leqslant N_{\ell}(n,t).$
5. Conclusion and Open Problems
In this paper we derive the maximum size of the intersection of the $t$-deletion balls centered at $x_{1},x_{2},\cdots,x_{\ell}$ where $x_{i}\in\{0,1\}^{n}$ for $i\in\{1,2,\cdots,j\}$ and $x_{i}\neq x_{j}$ for $i\neq j$. Therefore, the minimum number of the deletion channel outputs is $N_{\ell}(n,t)+1$ in order to reconstruct a list of size $\ell-1$ of candidate sequences, one of which corresponds to the original transmitted sequence.
Since our results concern the binary case, the first open question is the following.
Open Question 1. What is the maximum possible size of the intersection of the $t$-deletion balls centered at $x_{1},x_{2},\dots,x_{\ell}$, where $x_{i}\in\{0,1,\dots,q-1\}^{n}$ for all $i\in\{1,2,\dots,\ell\}$ and $x_{i}\neq x_{j}$ for $i\neq j$?
Since the only condition above is $x_{i}\neq x_{j}$ for $i\neq j$, the following naturally arises as a second open question.
Open Question 2. What is the maximum possible size of the intersection of the $t$-deletion balls centered at $x_{1},x_{2},\dots,x_{\ell}$, where $x_{i}\in\{0,1,\dots,q-1\}^{n}$ for all $i\in\{1,2,\dots,\ell\}$ and $d_{L}(x_{i},x_{j})\geqslant d$ for $i\neq j$?
Acknowledgment
The author would like to thank Prof. Alexander Barg for suggesting this problem and providing valuable guidance. Additionally, the author extends thanks to Yihan Zhang for insightful discussions.